1. A 560.0 N cyclist sits on a bicycle seat and compresses the two springs that support it.  The spring constant equals 2.2 X 104 N/m. 
  2. How much is each spring compressed?
  3. By how much does the compression increase in each spring’s elastic PE?
1
Expert's answer
2021-05-07T09:45:00-0400

1.How much is each spring compressed?

"F =kx"

where x is the compression. Since there are two strings, the force will be distributed on them

"F =2kx"

"\displaystyle x = \frac{F}{2k} = \frac{560}{2 \cdot2.2 \cdot 10^4}= 12.73 \cdot 10^{-3}\, m = 12.73\; mm"

Each string is compressed by 12.73 mm.


2.By how much does the compression increase in each spring’s elastic PE?

The increase in each spring’s elastic potential energy is

"\displaystyle E = \frac{kx^2}{2}= \frac{2.2 \cdot 10^4 \cdot 12.73^2 \cdot 10^{-6}}{2} = 178.26 \cdot 10^{-2} = 1.78 \; J"

PE of each spring increases by 1.78 J.